A 0.250-kg lump of clay is dropped from a height oL10.7
Answer #1

The impulse is equal to the decrease of the momentum of the lump of clay. As the clay falls 1.45 meters, its velocity increases at the rate of 9.8 m/s each second. Let’s use the following equation to determine its velocity when it hits the floor.

vf^2 = vi^2 + 2 * a * d, vi = 0
vf^2 = 2 * 9.8 * 1.45
vf = √28.42
This is approximately 5.33 m/s.

Momentum = 0.250 * 5.33

This is approximately 1.33 kg * m/s. As the floor exerts an upward force on the lump of clay, its velocity will decrease to 0 m/s. My point is the impulse is equal to the momentum when it hit floor.

Impulse = Force * time
9.5 ms is the total time the force is applied. If we use this time, we will get the average force. The maximum force will be two times the average force.
t = 0.0095 seconds.

F * 0.0095 = 0.250 * 5.33
F = 0.250 * 5.33/0.0095 = 140.3
This is approximately140 N.
Maximum force = 280 N